(c) Show directly that iC-Ax, then C(Dx) Dx) 7. Consider any A and a Givens rotation M in the 1-2 plane Cl cos θ -sinθ 0 A de fl , M |sin θ cos θ 0 Choose the rotation angle to produce zero in the (3, 1) entry of M- AM 336 Chapter 5 Eigenvalues and Eigenvectors Note. This zeroing is not so easy to continue, because the rotations that produce zero in place of d and h will spoil the new zero in the corner. We have to leave one diagonal below the main one, and finish the eigenvalue calculation in a different way Otherwise, if we could make A diagonal and see its eigenvalues, we would be finding the roots of the polynomial det(A -XI) by using only the square roots that determine cos θ-and that is impossible 8. What matrix M changes the basis V-(1,1), V2 (1,4) to the basis (2,5), v2 = (1,4)? The columns of M come from expressing Vi and V2 as combinations 2mijvi of the vs.
Answer #1

1) Given matrix is : A = ad g and M = \begin{bmatrix} \cos\theta & -\sin\theta & 0\\ \sin\theta & \cos\theta & 0\\ 0 & 0 & 1 \end{bmatrix} .

Now, M-1 = MT   [Since here M is an orthogonal matrix]

i.e., M-1 = \begin{bmatrix} \cos\theta & \sin\theta & 0\\ -\sin\theta & \cos\theta & 0\\ 0 & 0 & 1 \end{bmatrix} .

Then, M-1AM = \begin{bmatrix} \cos\theta & \sin\theta & 0\\ -\sin\theta & \cos\theta & 0\\ 0 & 0 & 1 \end{bmatrix} \begin{bmatrix} a & b & c\\ d & e & f\\ g & h & i \end{bmatrix}\begin{bmatrix} \cos\theta & -\sin\theta & 0\\ \sin\theta & \cos\theta & 0\\ 0 & 0 & 1 \end{bmatrix}

We need the (3,1) entry of M-1AM. So we are not going to calculate three matrices. For (3,1) entry, we just calculate \begin{bmatrix} 0 & 0 & 1 \end{bmatrix} \begin{bmatrix} a & b & c\\ d & e & f\\ g & h & i \end{bmatrix}\begin{bmatrix} \cos\theta\\ \sin\theta\\ 0\end{bmatrix}.

Now, \begin{bmatrix} 0 & 0 & 1 \end{bmatrix} \begin{bmatrix} a & b & c\\ d & e & f\\ g & h & i \end{bmatrix}\begin{bmatrix} \cos\theta\\ \sin\theta\\ 0\end{bmatrix} = \begin{bmatrix} g & h & i \end{bmatrix} \begin{bmatrix} \cos\theta\\ \sin\theta\\ 0\end{bmatrix} = g\cos\theta+h\sin\theta

If both of g and h are non-zero, then (3,1) entry cannot be zero.

If h = 0 and g \neq 0, then the entry becomes g\cos\theta and it will be 0 if \theta=\frac{n\pi}{2} , where n is odd.

If g = 0 and h \neq =0, then the entry becomes h\sin\theta and it will be 0 if \theta=\frac{n\pi}{2} , where n is even.

If h = 0 and g = 0, then the entry becomes 0 and it does not depend on \theta .

2) Given, V1 = (1,1), V2 = (1,4) and v1 = (2,5), v2 = (1,4).

Now, V1 = (1,1) = 1*(2,5) + (-1)*(1,4) = 1*v1+(-1)*v2

And, V2 = (1,4) = 0*(2,5) + 1*(1,4) = 0*v1+1*v2

Therefore, the required matrix is : M = \begin{bmatrix} 1 & 0\\ -1 & 1 \end{bmatrix} .